(x/3-y/2)(x/3-y/2)=
Answer by checkley75(3666)

(x/3-y/2)(x/3-y/2) using the FOIL process we get:
x/3*x/3=x^2/9
x/3*-y/2=-xy/6
-y/2*x/3=-xy/6
-y/2*-y/2=y^2/4 now add/combine these terms
x^2/9-2xy/6+y^2/4
x^2/9-xy/3+y^2/4